Assume an economy can be represented by Y = 10 + 0.8×C + G, where Y is income (GDP), G is government spending and C is consumption. If G was to rise from 100 to 200 by how much would the equilibrium value of Y increase?
a. 500 b. 80 c. 100 d. 200
1
Expert's answer
2016-06-03T09:39:02-0400
Y = 10 + 0.8×C + G.
If G was to rise from 100 to 200, then the change of equilibrium value of Y will be:
Y1 = 10 + 0.8*C + 100 = 0.8C + 110,
Y2 = 10 + 0.8*C + 200 = 0.8C + 210,
Y2 - Y1 = 100, so the correct answer is c.